# CLIP-MoE: Towards Building Mixture of Experts for CLIP with Diversified Multiplet Upcycling

### v1

@@ Abstract
In recent years, Contrastive Language-Image Pre-training (CLIP) has become a cornerstone in multimodal intelligence. However, recent studies have identified that the information loss in the CLIP encoding process is substantial, and CLIP tends to capture only coarse-grained features from the input. This deficiency significantly limits the ability of a single CLIP model to handle images rich in visual detail. In this work, we propose a simple yet effective model-agnostic strategy, Diversified Multiplet Upcycling (DMU), for CLIP. This strategy innovatively mitigates information loss in CLIP multimodal alignment by integrating multiple CLIP models that capture diversified, complementary information into a CLIP Mixture of Experts (MoE) with sparse activation. However, achieving such multiplet CLIP models is far from straightforward, as existing open-source CLIP models cannot directly form homogeneous experts or capture diversified complementary information. Inspired by the recently proposed Multistage Contrastive Learning (MCL), which constructs multiple CLIP models that share the same structure while capturing different complementary information, Diversified Multiplet Upcycling efficiently fine-tunes a series of CLIP models that capture different feature spaces, from a dense pre-trained CLIP checkpoint, sharing parameters except for the Feed-Forward Network (FFN). These models can then be transformed into a CLIP-MoE with a larger model capacity, leading to significantly enhanced performance with minimal computational overhead. To the best of our knowledge, Diversified Multiplet Upcycling is the first approach to introduce sparsely activated MoE into CLIP foundation models, whereas previous methods have focused solely on vision representations or model-wise ensembling. Extensive experiments demonstrate the significant performance of CLIP-MoE across various zero-shot retrieval, zero-shot image classification tasks, and downstream Multimodal Large Language Model (MLLM) benchmarks by serving as a vision encoder. Furthermore, Diversified Multiplet Upcycling enables the conversion of any dense CLIP model into CLIP-MoEs, which can seamlessly replace CLIP in a plug-and-play manner without requiring further adaptation in downstream frameworks. Through Diversified Multiplet Upcycling, we aim to provide valuable insights for future research on developing more efficient and effective multimodal learning systems. Codes and models are released at: https://github.com/OpenSparseLLMs/CLIP-MoE.
@@ 1 Introduction
To address the above limitations, we propose a simple yet effective model-agnostic strategy, Diversified Multiplet Upcycling (DMU), for CLIP, which leverages the sparsely activated Mixture of Experts (MoE) framework to extend model capacity while fully utilizing off-the-shelf pre-trained dense checkpoints, avoiding the need for training from scratch. MoE has proven effective in scaling large pre-trained models by using fixed activated parameters, enhancing both performance and robustness (Jiang et al., 2024; Dai et al., 2024; Chen et al., 2024a). In Diversified Multiplet Upcycling, we first fine-tune the base dense CLIP model to produce a series of multiplet CLIP models using the recently proposed Multistage Contrastive Learning (MCL) (Zhang et al., 2024b). MCL generates models that encode diversified information through a multistage clustering and fine-tuning process. By multiplet, we refer to CLIP models that share all parameters except for the feed-forward network (FFN) layers during MCL fine-tuning. By diversified, we mean that these models yield a series of FFN experts, each capturing different aspects of the input information, which are then used to initialize a CLIP-MoE model. Finally, through continuous fine-tuning of the router in CLIP-MoE, we ensure the full utilization of all experts, enabling CLIP-MoE to capture richer and more useful information than the base model, while leveraging sparse activation to avoid the explosion of activated parameters.
@@ 3.1 Multistage Contrastive Learning (MCL)
Formally, let $\bm{X}=\{\mathbf{x}_{i}\}_{i=1}^{M}$ represent a dataset, where $x_{i}$ denotes …. After training the encoder in the first stage, we obtain encoded representations $\bm{Z}_{0}=\{f_{0}(\mathbf{x}_{i})\}_{i=1}^{M}$. By clustering $\bm{Z}_{0}$, we obtain cluster assignments $\bm{Y}_{0}=\{\mathbf{y}_{(i,0)}\}_{i=1}^{M}$. In the $j^{th}$ stage, after the cluster-and-contrastive process, each sample $\mathbf{x}_{i}$ is assigned to an accumulated cluster $\hat{\mathbf{y}}_{(i,j)}=[\mathbf{y}_{(i,0)},\cdots,\mathbf{y}_{(i,j-1)}]$. The objective at the $j^{th}$ stage is to optimize:
@@ 3.2 Mixture of Experts (MoE)
where $R(\mathbf{x})_{i}$ denotes the $i$-th component of the routing weight vector produced by the router network $\mathbf{W}_{r}$.
@@ 4.3 Continuous Fine-Tuning of CLIP-MoE
Following Fedus et al. (2022b), we set $\alpha=0.01$ by default. By applying MoE-Packing to CLIP, we obtain a CLIP-MoE model that is capable of capturing more useful information than the base model, with minimal computational overhead, resulting in a robust and efficient enhancement of the CLIP model.
@@ 5.1 Datasets
ShareGPT4V. ShareGPT4V (Chen et al., 2023) is a high-quality image-text dataset containing 1.2 million highly descriptive captions. The captions are generated by a Multimodal Large Language Model (MLLM) fine-tuned on 100k image-text pairs produced by GPT4V, resulting in well-aligned image-text pairs. In this paper, we xx
@@ 5.2 Baselines
Direct Fine-tuning. As our experiments incorporate additional data, we use direct fine-tuning as a basic baseline to serve as a reference for evaluating the performance contributions from the additional data.
@@
Long-CLIP. Long-CLIP (Zhang et al., 2024a) introduces an efficient pipeline to enhance CLIP performance through fine-tuning on high-quality image-caption datasets with long captions. It aligns the long caption of an image with the encoded image features and aligns the short caption with the primary components of the image features. While effective on the ShareGPT4V dataset, Long-CLIP is limited to datasets with a similar structure, where each image has both one short and one long caption. Moreover, it requires significantly more computational resources compared to our approach.
@@ 5.5 Evaluation
Following the methodology outlined in Zhang et al. (2024a), we evaluate text-to-image (T2I) and image-to-text (I2T) retrieval on the 5k COCO validation set (Lin et al., 2014) and the 30k Flickr30k (Young et al., 2014) dataset. The results are presented in Table 1. Given that both Recap-DataComp-1M and ShareGPT4V datasets offer higher caption quality and longer average caption lengths compared to web datasets, Direct Fine-Tuning, Sparse Upcycling, and CLIP-MoE demonstrate superior performance over the original OpenAI model across most tasks, including COCO I2T, COCO T2I, and Flickr T2I. However, for Flickr I2T, Sparse Upcycling, and Direct Fine-Tuning show significant performance degradation on the Recap-DC dataset. In this fine-tuning context, Sparse Upcycling only provides a limited advantage over Direct Fine-Tuning. Although Long-CLIP clearly outperforms both Direct Fine-Tuning and Sparse Upcycling, it is incompatible with the Recap-DataComp dataset, because it requires each image to have both a short and a long caption. In contrast, our proposed CLIP-MoE surpasses all baselines on both Recap-DataComp and ShareGPT4V, maintaining consistent performance without degradation by leveraging the diverse information extracted by MoE experts initialized through different stages of MCL.
@@ 5.6 Discussion
Figure 2: Proportion of tokens assigned to each expert on the COCO validation dataset and ImageNet validation dataset. Here, we consider experts that are either selected as a first or second choice by the router.
@@ 6 Conclusion & Future Work
In this paper, we proposed a novel Diversified Multiplet Upcycling for CLIP to enhance the model with minimal computational overhead. Our method enables the extraction of diversified and complementary experts across multiple fine-tuning stages, which are then utilized within the MoE framework to capture richer information from the inputs. This approach is straightforward to apply, model-agnostic, and provides a new path to scale and improve CLIP foundation models. By leveraging off-the-shelf CLIP checkpoints and newly constructed high-quality image-text datasets, our method avoids the costly process of training CLIP models from scratch. We demonstrated the effectiveness and efficiency of our approach through extensive experiments across various datasets and tasks.
For future work, our current experiments are limited to image and text modalities. We plan to extend our method to additional modalities, such as audio and video. Beyond the fine-tuning settings explored in this paper, we aim to experiment with larger datasets and test large-scale continuous training settings to further explore the scalability and performance boundaries of Diversified Multiplet Upcycling. Additionally, while we tested CLIP-MoE as a vision encoder for MLLMs, we will also investigate its potential as a text encoder in generative tasks, such as in stable diffusion.

### v2

@@ Abstract
In recent years, Contrastive Language-Image Pre-training (CLIP) has become a cornerstone in multimodal intelligence. However, recent studies have identified that the information loss in the encoding process of CLIP is substantial. Such deficiency significantly limits the ability of a single CLIP model to handle images rich in visual detail. In this work, we propose a simple yet effective model-agnostic strategy, Diversified Multiplet Upcycling (DMU) for CLIP. It integrates multiple CLIP models that capture diversified, complementary information into a Mixture of Experts (MoE) architecture. Inspired by the recently proposed Multistage Contrastive Learning (MCL), which constructs multiple CLIP models that share the same structure while capturing different complementary information, Diversified Multiplet Upcycling efficiently fine-tunes a series of CLIP models from a dense pre-trained CLIP checkpoint to capture different feature distributions, sharing parameters except for the Feed-Forward Network (FFN). These models are then transformed into a CLIP-MoE with a larger model capacity but minimal computational overhead. Extensive experiments demonstrate the significant performance of CLIP-MoE across various zero-shot retrieval, zero-shot image classification tasks, and downstream Multimodal Large Language Model (MLLM) benchmarks by serving as a vision encoder. Furthermore, Diversified Multiplet Upcycling enables the conversion of any dense CLIP model into CLIP-MoEs, which can seamlessly replace CLIP in a plug-and-play manner without requiring further adaptation in downstream frameworks. Through Diversified Multiplet Upcycling, we aim to provide valuable insights for future research on developing more efficient and effective multimodal learning systems. Codes and models are released at: https://github.com/OpenSparseLLMs/CLIP-MoE.
@@ 1 Introduction
To address the above limitations, we propose a simple yet effective model-agnostic strategy, Diversified Multiplet Upcycling (DMU), for CLIP, which leverages the sparsely activated Mixture of Experts (MoE) framework to extend model capacity while fully utilizing off-the-shelf pre-trained dense checkpoints, avoiding the need for training from scratch. MoE has proven effective in scaling large pre-trained models by using fixed activated parameters, enhancing both performance and robustness (Jiang et al., 2024; Dai et al., 2024; Chen et al., 2024a). In Diversified Multiplet Upcycling, we first fine-tune the base dense CLIP model to produce a series of multiplet CLIP models using the recently proposed Multistage Contrastive Learning (MCL) (Zhang et al., 2024b). MCL generates models that encode diversified information through a multistage clustering and fine-tuning process. By multiplet, we refer to CLIP models that share all parameters except for the feed-forward network (FFN) layers during MCL fine-tuning. By diversified, we mean that these models yield a series of FFN experts, each capturing different aspects of the input information, which are then used to initialize a CLIP-MoE model. Finally, through fine-tuning the router in CLIP-MoE, we ensure the full utilization of all experts, enabling CLIP-MoE to capture richer and more useful information than the base model, while leveraging sparse activation to avoid the explosion of activated parameters.
@@ 3.1 Multistage Contrastive Learning (MCL)
Formally, let $\bm{X}=\{\mathbf{x}_{i}\}_{i=1}^{M}$ represent a dataset. After training the encoder in the first stage, we obtain encoded representations $\bm{Z}_{0}=\{f_{0}(\mathbf{x}_{i})\}_{i=1}^{M}$. By clustering $\bm{Z}_{0}$, we obtain cluster assignments $\bm{Y}_{0}=\{\mathbf{y}_{(i,0)}\}_{i=1}^{M}$. In the $j^{th}$ stage, after the cluster-and-contrastive process, each sample $\mathbf{x}_{i}$ is assigned to an accumulated cluster $\hat{\mathbf{y}}_{(i,j)}=[\mathbf{y}_{(i,0)},\cdots,\mathbf{y}_{(i,j-1)}]$. The objective at the $j^{th}$ stage is:
@@ 3.2 Mixture of Experts (MoE)
where $R(\mathbf{x})_{i}$ denotes the $i$-th routing weight vector produced by the router network $\mathbf{W}_{r}$.
@@ 4.3 Continuous Fine-Tuning of CLIP-MoE
Following Fedus et al. (2022b), we set $\alpha=0.01$ by default. By applying MoE-Packing to CLIP, we obtain a CLIP-MoE model that is capable of capturing more useful information than the base model, with minimal computational overhead, resulting in a robust and efficient enhancement of CLIP.
@@ 5.1 Datasets
ShareGPT4V. ShareGPT4V (Chen et al., 2023) is a high-quality image-text dataset containing 1.2 million highly descriptive captions. The captions are generated by a Multimodal Large Language Model (MLLM) fine-tuned on 100k image-text pairs produced by GPT4V, resulting in well-aligned image-text pairs.
@@ 5.2 Baselines
Direct Fine-tuning. As our experiments incorporate additional data, we use direct fine-tuning as a basic baseline to evaluate the performance contributions from the additional data.
@@
Long-CLIP. Long-CLIP (Zhang et al., 2024a) introduces an efficient pipeline to enhance CLIP performance through fine-tuning on high-quality image-caption datasets with long captions. It aligns the long caption of an image with the encoded image features and the short caption with the primary components of the image features. While effective on the ShareGPT4V dataset, Long-CLIP is limited to datasets with a similar structure, where each image has both one short and one long caption. Moreover, it requires significantly more computational resources compared to our approach.
@@ 5.5 Evaluation
Following the methodology outlined in Zhang et al. (2024a), we evaluate text-to-image (T2I) and image-to-text (I2T) retrieval on the 5k COCO validation set (Lin et al., 2014) and the 30k Flickr30k (Young et al., 2014) dataset. The results are presented in Table 1. Given that both Recap-DataComp-1M and ShareGPT4V datasets offer higher caption quality and longer average caption lengths compared to web datasets, Direct Fine-Tuning, Sparse Upcycling, and CLIP-MoE demonstrate superior performance over the original OpenAI model across most tasks, including COCO I2T, COCO T2I, and Flickr T2I. However, for Flickr I2T, Sparse Upcycling, and Direct Fine-Tuning show significant performance degradation on the Recap-DC dataset. In this fine-tuning context, Sparse Upcycling only provides a limited advantage over Direct Fine-Tuning. Although Long-CLIP clearly outperforms both Direct Fine-Tuning and Sparse Upcycling, it is incompatible with the Recap-DataComp dataset, because it requires each image to have both a short and a long caption. In contrast, our proposed CLIP-MoE surpasses all baselines on both Recap-DataComp and ShareGPT4V, maintaining consistent performance by leveraging the diverse information extracted by MoE experts initialized through different stages of MCL.
@@ 5.6 Discussion
Figure 2: Proportion of tokens assigned to each expert on the COCO and ImageNet validation dataset. Here, we consider experts that are either selected as a first or second choice by the router.
@@ 6 Conclusion & Future Work
In this paper, we proposed a novel Diversified Multiplet Upcycling for CLIP to enhance the model with minimal computational overhead. Our method enables the extraction of diversified and complementary experts across multiple fine-tuning stages, which are then utilized within the MoE framework to capture richer information from the inputs. This approach is straightforward to apply, model-agnostic, and provides a new path to scale and improve CLIP foundation models. By leveraging off-the-shelf CLIP checkpoints and newly constructed high-quality image-text datasets, our method avoids the costly process of training CLIP models from scratch. For future work, our current experiments are limited to image and text modalities. We plan to extend our method to additional modalities, such as audio and video. Beyond the fine-tuning settings explored in this paper, we aim to experiment with larger datasets and test large-scale continuous training settings to further explore the boundaries of Diversified Multiplet Upcycling. Additionally, while we tested CLIP-MoE as a vision encoder for MLLMs, we will also investigate its potential as a text encoder in generative tasks, such as in stable diffusion.